OpenStudy (anonymous):

How many solutions will this system of equations have? x - 2y = 24 3x - 6y = 72 No solution Infinite solutions One solution Two solutions

OpenStudy (anonymous):

second eq. is same except that it is multiplied by 3 so actually there is only one eq. and it has infinite solutions.

OpenStudy (anonymous):

@surjithayer is right they are both the same relation, so whatever set of solutions works in the one also works in the other

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x - 2y = 24 3x - 6y = 72 (divide through by 3) --> x - 2y = 24 they are the same equations therefore share an infinite number of solutions infinite number of solutions is for when system of equations is the same equations no solution is when they are parallel (slopes are the same)